# AIMC-Spec: A Benchmark Dataset for Automatic Intrapulse Modulation Classification under Variable Noise Conditions

### v1

@@ VI Conclusion
This paper presents AIMC-Spec, a synthetic dataset purpose-built for AIMC using spectrogram-based image classification techniques. It includes a diverse range of 33 intrapulse modulation types—spanning frequency, phase, and hybrid categories—across a wide SNR spectrum, allowing for rigorous and structured benchmarking under varied signal conditions.
Five representative deep learning algorithms were implemented and evaluated on AIMC-Spec, highlighting differing behaviors across model architectures. The results are consistent with known trends in AIMC: frequency modulations are generally more separable, while phase and hybrid types present greater challenges. The benchmarking demonstrates AIMC-Spec’s potential as a diagnostic tool for analyzing algorithm performance across modulation types and noise levels.
@@ VI-A Future Works
Future work should extend AIMC-Spec to include additional modulation classes and finer grained SNR intervals to better assess model robustness in harsh conditions. Furthermore, more benchmarking algorithms, including those not originally designed for AIMC, should be explored to evaluate cross-domain generalizability.
@@
Beyond architectural benchmarking, future iterations of AIMC-Spec will incorporate greater environmental realism by integrating measured or parameter-derived radar signals and by simulation operation effects such as multipath propagation, fading, pulse dropouts, phase jitter, and measurement noise. These additions will enable the dataset to emulate real-world conditions more accurately and provide a stronger basis for evaluating resilience and generalization performance.
With ongoing updates, AIMC-Spec can evolve into a lasting benchmark for the AIMC community. Its structured, transparent design enables reproducible research, fair comparisons, and clearer insights into model behavior-providing a foundation for future research in intrapulse modulation classification.

### v2

@@ VI Conclusion
This paper presents AIMC-Spec, a synthetic dataset designed to support benchmarking of AIMC using spectrogram-based image classification techniques. The dataset includes 30 intrapulse modulation types spanning frequency and phase categories, and evaluates performance across a controlled SNR range from –6 dB to +6 dB. By combining diverse modulation classes with realistic noise conditions, AIMC-Spec provides a structured environment for assessing model robustness under varying signal quality.
Five representative deep learning algorithms were implemented and evaluated on AIMC-Spec, including LDC-Unet, ViT, LPI-Net, CDAE-DCNN, and STFT-CNN. The benchmarking experiments revealed clear differences in performance across model architectures, with LDC-Unet consistently achieving the highest accuracy across both the full modulation task and the FM subset. CNN-based architectures with dense connections and multi-scale feature extraction demonstrated greater resilience to noise, while transformer and lightweight CNN approaches showed comparatively reduced robustness.
Overall, the experimental results highlight the complexity of the AIMC-Spec classification task, particularly when distinguishing phase and hybrid modulations whose spectral characteristics overlap under noisy conditions. These findings demonstrate that AIMC-Spec can serve not only as a benchmarking dataset but also as a diagnostic tool for understanding how different model architectures respond to challenging signal environments.
@@ VI-A Future Works
Future work will extend AIMC-Spec by incorporating additional intrapulse modulation types and expanding the diversity of signal conditions used during dataset generation. Including a broader range of modulation behaviors will allow more comprehensive benchmarking and help evaluate model performance across increasingly complex classification tasks.
@@
Beyond architectural benchmarking, future iterations of AIMC-Spec will incorporate greater environmental realism by integrating measured or parameter-derived radar signals and by simulating operational effects such as multipath propagation, fading, pulse dropouts, phase jitter, and measurement noise. These additions will enable the dataset to emulate real-world conditions more accurately and provide a stronger basis for evaluating algorithm robustness and generalization performance.
With continued development, AIMC-Spec can evolve into a comprehensive benchmark for the AIMC community. Its structured and transparent design supports reproducible experimentation, fair comparison between algorithms, and deeper analysis of how model architectures respond to challenging signal environments.
Appendix A Compressed Benchmark Architectures
Figure 5: Compressed architecture of CDAE-DCNN.
Figure 6: Compressed architecture of LDC-Unet.
Figure 7: Compressed architecture of LPI-Net.
Figure 8: Compressed architecture of STFT-CNN.
Figure 9: Compressed architecture of ViT.
Appendix B Tabular Format of Graphs
TABLE IV: Accuracy (%) vs SNR (dB) for the full modulation classification task. Best values per row are in bold.
SNR STFT-CNN LPI-Net CDAE-DCNN LDC-Unet ViT
6-6 24.06 47.47 42.59 57.37 49.07
3-3 31.18 53.44 49.06 61.85 55.19
0 32.74 64.91 57.42 71.45 66.38
+3+3 37.58 64.59 57.55 70.04 65.57
+6+6 38.68 69.43 61.13 74.29 70.42
TABLE V: Accuracy (%) vs SNR (dB) for the frequency-modulation-only classification task. Best values per row are in bold.
SNR STFT-CNN LPI-Net CDAE-DCNN LDC-Unet ViT
6-6 51.56 75.67 72.21 81.24 76.54
3-3 64.71 81.07 78.21 86.24 82.35
0 67.41 85.47 81.17 90.12 86.61
+3+3 73.31 86.17 82.04 91.02 87.32
+6+6 73.01 87.47 81.91 91.30 88.14
Appendix C List of Acronyms
TABLE VI: Intrapulse Modulation Types and Descriptions
Acronym AIMC-Spec Full Name Description
UNMOD Unmodulated Constant frequency, no modulation across the pulse duration.
LFM Linear Frequency Modulation Frequency increases (LFM_up) or decreases (LFM_down) linearly across time (chirp).
NLFM Nonlinear Frequency Modulation Frequency changes non-linearly to reduce sidelobes.
SFM Step Frequency Modulation Discrete frequency steps within the pulse.
EQFM Equal-Width Frequency Modulation Uniform sub-bands of frequency modulation.
DLFM Discrete Linear FM Linear FM in discrete blocks, can be either up_down or down_up.
MLFM Modified Linear FM Linear frequency modulation with nonlinear terms added to reduce sidelobes and enhance range resolution.
BFSK Binary Frequency Shift Keying Equivalent to 2FSK; binary frequency hopping.
4FSK 4-level Frequency Shift Keying Switches between four frequencies.
FSK Frequency Shift Keying General category of modulation using frequency switching.
SINFM Sinusoidal Frequency Modulation Frequency follows a sinusoidal curve across time.
TRIFM Triangular Frequency Modulation Symmetric linear up/down chirp within a pulse. AIMC-Spec considers DLFM_Up_Down as TriFM due to the single pulse classification task.
SCR Stepped Chirp Sequence of chirps with stepped frequency offsets.
PCR3 Piecewise Constant Ramp (3 segments) Frequency ramps in fixed intervals across 3 segments.
QFSK Quadrature Frequency Shift Keying Four FSK tones placed orthogonally.
BPSK Binary Phase Shift Keying Phase alternates between two states (0, $\pi$).
QPSK Quadrature Phase Shift Keying Phase modulated using four discrete values.
FRANK Frank Code Polyphase code with linear phase progression.
BARKER Barker Code Binary code with good autocorrelation properties. Consists of: 2_1, 2_2, 3, 4_1, 4_2, 5, 7, 11, 13
COSTAS Costas Code Frequency-hopping sequence with good ambiguity properties.
P1–4 P1–P4 Polyphase Codes Four distinct polyphase modulation schemes designed for pulse compression, each offering specific autocorrelation properties and sidelobe control for improved range resolution.
EXP Exponential FM Frequency varies exponentially over time.
T1-4 T1–T4 Codes Polyphase codes with tailored phase sequences for low sidelobes and Doppler tolerance, typically 16–100 chips long, similar to P1–P4 codes but optimized for Doppler performance.
LFM-BPSK LFM with BPSK LFM chirp modulated with BPSK phase switching.
2FSK-BPSK 2FSK with BPSK 2FSK combined with binary phase modulation.
LFM-SFM LFM with SFM Hybrid of linear chirp and step frequency.
BPSK-LFM BPSK with LFM BPSK applied over an LFM waveform.
EQFM-BPSK EQFM with BPSK Phase modulated EQFM waveform.
EQFM-2FSK EQFM with 2FSK Frequency hopping within EQFM envelope.
FRANK-EQFM Frank-coded EQFM EQFM waveform modulated using Frank coding.
FRANK-4FSK Frank-coded 4FSK 4FSK with Frank phase coding.
LFM-4FSK LFM and 4FSK hybrid Combines LFM’s linear chirp with 4FSK’s four discrete frequency steps to improve range resolution and interference rejection.
BPSK-FSK BPSK with FSK Binary phase shift applied to FSK waveform.
FSK-LFM FSK with LFM Frequency-hopping waveform with LFM characteristics.
FSK-BPSK FSK with BPSK Frequency shift keyed waveform with binary phase.